OpenStudy (anonymous):

The variables x and y vary inversely. If x=13/5 and y=5, write an equation that relates to x and y. please solve

OpenStudy (zzr0ck3r):

y=k/x plug in y = 5, x = 13/5 5 = k/(13/5) = k(5/13) k = 13

OpenStudy (zzr0ck3r):

understand?

OpenStudy (anonymous):

kinda

OpenStudy (zzr0ck3r):

can you give me the equation with slope 2, passing through the point (2,3)?

OpenStudy (zzr0ck3r):

donnt actually do it, but could you?